# Urban highways are barriers to social ties

### v1

@@ Abstract
Cities are hubs of concentrated social capital that can foster diversity and innovation (1, 2). However, this potential is threatened by spatial fragmentation through built infrastructure that can separate neighborhoods (3, 4), exacerbate inequalities (5, 6), and contribute to segregation (7). Among various types of barriers fragmenting urban areas, roads designed for motorized traffic are the most ubiquitous, especially highways (8, 4, 14). Since the 1960s, urban planners have theorized that high-traffic roads reduce opportunities for creating and maintaining social ties across divided neighborhoods (10), thus undermining the social cohesion essential for the development of thriving communities. This premise lies at the core of contemporary urban planning research and interventions (11, 12) that strive to meet the UN’s sustainable development goal of “making cities and human settlements inclusive, safe, resilient and sustainable” (13).
@@
Notoriously, urban highways in the US have been instrumentalized for government-backed racial segregation, creating social divides between communities that persist to this day (14, 27). We therefore revisit several highways in US cities that are well-documented for their historic role in racial segregation, finding potential evidence for long-lasting effects several decades after their construction, by measuring high Barrier Scores in contemporary social networks.
@@ Results
To highlight the practical implications of our quantitative findings, we now frame them within a broader historical context, with a particular focus on racial residential segregation. Race is only one of many social categories that can influence the formation of social connections. However, the Interstate Highway System – which we study here in the urban context – is highly relevant for aggravating racial segregation in US cities (79), making the association of our Barrier Score with racial residential segregation a compelling case study. Overwhelming historical records show how urban highway construction in the name of “urban renewal” has been frequently used as a racist policy toolbox to purposefully disrupt or isolate Black neighborhoods (33), together with other de jure segregation tools like redlining and housing policy (14). Such exclusionary urban policies, put in place decades ago, have literally cemented racial divides in US cities and have therefore not lost any of their societal relevance today (15, 27). Indeed, the US Department of Transportation acknowledges the issue in its 2023 “Reconnecting Communities Pilot Program”, an “initiative to reconnect communities that are cut off from opportunity and burdened by past transportation infrastructure decisions” (35).
@@
Next, we discuss the three cities with highways as interracial barriers. In Oklahoma City (Fig. 4D), the “urban renewal” highway construction projects had particularly dire impacts on historically Black neighborhoods such as Deep Deuce (51). As of today, the I-235 in Oklahoma City remains a clearly perceived division line between majority Black and majority white neighborhoods (49). In Detroit (Fig. 4E), the construction of several highways during “urban renewal” erased and eroded numerous historically Black neighborhoods such as Black Bottom and Paradise Valley (29). Here, “expressway displacement” (29) combined with pronounced discrimination led to several housing crises over the last decades, severely impacting the Black population. Lastly, in the city of Austin (Fig. 4F), the I-35 was built along East Avenue, an intentionally enforced segregation line whose impacts are visible up to this day (65). At the same time, the I-35, for which expansion plans are currently underway with 4 billion USD allocated (70, 27), stands out with a high Barrier Score.
Finally, three cities from our case studies are well-known for their intraracial highway barriers. Columbus (Fig. 4G) is a particularly startling example of highway construction as deliberate neighborhood destruction (74), with today’s highway routes aligning with former redlining maps. The most severely impacted neighborhoods like Flytown, Hanford Village, or Bronzeville, were economically disadvantaged and predominantly Black; at the same time, the closeby but predominantly white, affluent neighborhood Bexley was spared from the highways (73, 74). In Richmond (Fig. 4H), highway construction and segregationist housing policies interacted to create a “concentration of racialized poverty” (25) that lasts until the present day. Richmond’s neighborhood of Jackson Ward, formerly dubbed “Black Wall Street”, was bisected by the I-95 and the I-64/I-95 interchange, ultimately leading to its decline. Finally, in Nashville (Fig. 4I), the I-40 was routed through a bustling Black neighborhood without any appraisal of potential consequences for the community, bisecting the once-thriving Jefferson Street, and at a larger scale undermining Black commercial and educational institutions, decisively contributing to today’s high poverty rates in the area (78).
@@
To be clear, our approach is so far strictly correlational and cannot establish causality: from static data it is impossible to determine how thinned-out social ties across a highway section already were before its construction, say because of an existing racial divide (29, 14); or to which extent a new highway caused social ties to thin out. Scrutinizing causality would require longitudinal data, for example before and after the construction or removal of an urban highway. Nevertheless, within the historical context, our results paint a clear picture. Thus, our research could already help remediate previous political failures (16, 14) and enrich the debate on contemporary highway policies (70, 35, 27), to account for exclusionary effects of infrastructure, and to inform reparative justice approaches (55, 23). More generally, our research contributes to a more careful, evidence-based consideration of the social fabric in urban planning.
@@
We rely on an existing collection of geo-referenced tweets posted between 2012 and 2013, when the Twitter mobile app’s default setting was to annotate all tweets with the precise geographic coordinates at the time of posting. Previous work (1) used the friend-of-friend algorithm to identify the home locations of users with a sufficient number of posts with high accuracy. The dataset comes with the full network of mutual followership among all users whose home location is within the 50 most populous metropolitan areas in the United States. Overall, the network contains 982,459 users and 2,711,185 social ties between them. This dataset has proven to be a reliable resource to study spatial social networks within cities (4, 56). The home location estimation procedure, present statistics on the data, and its representativeness are described in detail in SI Sections 0.1 and 0.2.
@@ 0.12 The Color of Highways: The racial context of US highway construction
Social segregation can happen along many different axes; one of the most obvious ones to scrutinize in contemporary US cities is race. Segregationist practices were ruled out by law only in the 1960s, with the Civil Rights Acts of 1964 and 1968, the former outlawing “discrimination or segregation on the ground of race, color, religion, or national origin” (12) and the latter explicitly expanding this principle to the provision (selling and renting) of housing (13). Prior to this, however, explicitly racist and exclusionary urban policies such as redlining and racial zoning were widely practiced (14), and their legacy is still lingering up until today (15). The Interstate Highway System (IHS) is a compelling example thereof (16, 17). From the beginning of its construction in 1956, the IHS fostered urban sprawl for decades to come and played a major role in the suburbanization of US cities, which in turn is racially biased (18, 19, 20, 21). By the time that massive government-funded highway construction got underway throughout the country in the late 1950s, many US urban areas had majority Black inner cities. Many cities followed the advice of urban planner Robert Moses, not only on his proclaimed imperative that “most of our new (…) expressways (…) must go right through cities and not around them” (22), but also in his unambiguously racialized suggestion that “the practical solutions of the traffic problems in cities should be coordinated with slum clearance” (22), which meant building highways through Black neighborhoods. The choice of specific sites within a city where highways would lead through sparked a great number of protests across the country, known as Freeway Revolts (23). There are numerous, extensively studied examples of highway construction sites tearing apart or displacing historically Black communities (24, 25, 26, 27), proverbially known as “White roads through Black bedrooms” (28). Likewise, many decision-makers used such major construction projects as an instrument to get rid of so-called “ghettos” or to spatially exclude underprivileged communities (29, 14, 30). Effects of both practices are visible in today’s patterns of racial residential segregation. Below, we explore these effects and their historical context for our nine case study cities.

### v2

@@ Abstract
Cities are hubs of concentrated social capital that can foster diversity and innovation (1, 2). However, this potential is threatened by spatial fragmentation through built infrastructure that can separate neighborhoods (3, 4), exacerbate inequalities (5, 6), and contribute to segregation (7). Among various types of barriers fragmenting urban areas, roads designed for motorized traffic are the most ubiquitous, especially highways (8, 4, 9). Since the 1960s, urban planners have theorized that high-traffic roads reduce opportunities for creating and maintaining social ties across divided neighborhoods (10), thus undermining the social cohesion essential for the development of thriving communities. This premise lies at the core of contemporary urban planning research and interventions (11, 12) that strive to meet the UN’s sustainable development goal of “making cities and human settlements inclusive, safe, resilient and sustainable” (13).
@@
Notoriously, urban highways in the US have been instrumentalized for government-backed racial segregation, creating social divides between communities that persist to this day (9, 27). We therefore revisit several highways in US cities that are well-documented for their historic role in racial segregation, finding potential evidence for long-lasting effects several decades after their construction, by measuring high Barrier Scores in contemporary social networks.
@@ Results
To highlight the practical implications of our quantitative findings, we now frame them within a broader historical context, with a particular focus on racial residential segregation. Race is only one of many social categories that can influence the formation of social connections. However, the Interstate Highway System – which we study here in the urban context – is highly relevant for aggravating racial segregation in US cities (32), making the association of our Barrier Score with racial residential segregation a compelling case study. Overwhelming historical records show how urban highway construction in the name of “urban renewal” has been frequently used as a racist policy toolbox to purposefully disrupt or isolate Black neighborhoods (33), together with other de jure segregation tools like redlining and housing policy (9). Such exclusionary urban policies, put in place decades ago, have literally cemented racial divides in US cities and have therefore not lost any of their societal relevance today (34, 27). Indeed, the US Department of Transportation acknowledges the issue in its 2023 “Reconnecting Communities Pilot Program”, an “initiative to reconnect communities that are cut off from opportunity and burdened by past transportation infrastructure decisions” (35).
@@
Next, we discuss the three cities with highways as interracial barriers. In Oklahoma City (Fig. 4D), the “urban renewal” highway construction projects had particularly dire impacts on historically Black neighborhoods such as Deep Deuce (41). As of today, the I-235 in Oklahoma City remains a clearly perceived division line between majority Black and majority white neighborhoods (42). In Detroit (Fig. 4E), the construction of several highways during “urban renewal” erased and eroded numerous historically Black neighborhoods such as Black Bottom and Paradise Valley (43). Here, “expressway displacement” (43) combined with pronounced discrimination led to several housing crises over the last decades, severely impacting the Black population. Lastly, in the city of Austin (Fig. 4F), the I-35 was built along East Avenue, an intentionally enforced segregation line whose impacts are visible up to this day (44). At the same time, the I-35, for which expansion plans are currently underway with 4 billion USD allocated (45, 27), stands out with a high Barrier Score.
Finally, three cities from our case studies are well-known for their intraracial highway barriers. Columbus (Fig. 4G) is a particularly startling example of highway construction as deliberate neighborhood destruction (46), with today’s highway routes aligning with former redlining maps. The most severely impacted neighborhoods like Flytown, Hanford Village, or Bronzeville, were economically disadvantaged and predominantly Black; at the same time, the closeby but predominantly white, affluent neighborhood Bexley was spared from the highways (47, 46). In Richmond (Fig. 4H), highway construction and segregationist housing policies interacted to create a “concentration of racialized poverty” (48) that lasts until the present day. Richmond’s neighborhood of Jackson Ward, formerly dubbed “Black Wall Street”, was bisected by the I-95 and the I-64/I-95 interchange, ultimately leading to its decline. Finally, in Nashville (Fig. 4I), the I-40 was routed through a bustling Black neighborhood without any appraisal of potential consequences for the community, bisecting the once-thriving Jefferson Street, and at a larger scale undermining Black commercial and educational institutions, decisively contributing to today’s high poverty rates in the area (49).
@@
To be clear, our approach is so far strictly correlational and cannot establish causality: from static data it is impossible to determine how thinned-out social ties across a highway section already were before its construction, say because of an existing racial divide (43, 9); or to which extent a new highway caused social ties to thin out. Scrutinizing causality would require longitudinal data, for example before and after the construction or removal of an urban highway. Nevertheless, within the historical context, our results paint a clear picture. Thus, our research could already help remediate previous political failures (36, 9) and enrich the debate on contemporary highway policies (45, 35, 27), to account for exclusionary effects of infrastructure, and to inform reparative justice approaches (55, 23). More generally, our research contributes to a more careful, evidence-based consideration of the social fabric in urban planning.
@@
We rely on an existing collection of geo-referenced tweets posted between 2012 and 2013, when the Twitter mobile app’s default setting was to annotate all tweets with the precise geographic coordinates at the time of posting. Previous work (26) used the friend-of-friend algorithm to identify the home locations of users with a sufficient number of posts with high accuracy. The dataset comes with the full network of mutual followership among all users whose home location is within the 50 most populous metropolitan areas in the United States. Overall, the network contains 982,459 users and 2,711,185 social ties between them. This dataset has proven to be a reliable resource to study spatial social networks within cities (28, 56). The home location estimation procedure, present statistics on the data, and its representativeness are described in detail in SI Sections 0.1 and 0.2.
@@ 0.12 The Color of Highways: The racial context of US highway construction
Social segregation can happen along many different axes; one of the most obvious ones to scrutinize in contemporary US cities is race. Segregationist practices were ruled out by law only in the 1960s, with the Civil Rights Acts of 1964 and 1968, the former outlawing “discrimination or segregation on the ground of race, color, religion, or national origin” (12) and the latter explicitly expanding this principle to the provision (selling and renting) of housing (13). Prior to this, however, explicitly racist and exclusionary urban policies such as redlining and racial zoning were widely practiced (14), and their legacy is still lingering up until today (15). The Interstate Highway System (IHS) is a compelling example thereof (16, 17). From the beginning of its construction in 1956, the IHS fostered urban sprawl for decades to come and played a major role in the suburbanization of US cities, which in turn is racially biased (18, 19, 20, 21). By the time that massive government-funded highway construction got underway throughout the country in the late 1950s, many US urban areas had majority Black inner cities. Many cities followed the advice of urban planner Robert Moses, not only on his proclaimed imperative that “most of our new (…) expressways (…) must go right through cities and not around them” (22), but also in his unambiguously racialized suggestion that “the practical solutions of the traffic problems in cities should be coordinated with slum clearance” (22), which meant building highways through Black neighborhoods. The choice of specific sites within a city where highways would lead through sparked a great number of protests across the country, known as Freeway Revolts (23). There are numerous, extensively studied examples of highway construction sites tearing apart or displacing historically Black communities (24, 25, 26, 27), proverbially known as “White roads through Black bedrooms” (28). Likewise, many decision-makers used such major construction projects as an instrument to get rid of so-called “ghettos” or to spatially exclude underprivileged communities (29, 9, 30). Effects of both practices are visible in today’s patterns of racial residential segregation. Below, we explore these effects and their historical context for our nine case study cities.